OpenStudy (anonymous):

Find the sum of the geometric sequence –3, 15, –75, 375, … when there are 8 terms and select the correct answer below.

OpenStudy (anonymous):

what is r?

OpenStudy (anonymous):

What is what?

OpenStudy (anonymous):

the common ratio

OpenStudy (anonymous):

Is it 5?

OpenStudy (anonymous):

Or -5?

OpenStudy (anonymous):

-5 looks right

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle {\rm S}_{\rm n}=\frac{a_1\times \left(1-{\rm r}^{\rm n}\right)}{1-{\rm r}} }\) this is the formula, for the sum of a geometric sequence that starts from some \(a_1\) and has n terms.

OpenStudy (solomonzelman):

you need to find r, identify the \(a_1\) that you are given, and tell me the number of terms (which you plug in for n). plug everything into the formula, and calculate.

OpenStudy (anonymous):

I think I did something wrong in the equation because I got 24,414,063

OpenStudy (solomonzelman):

holly molly.....

OpenStudy (solomonzelman):

yeah it is wrong

OpenStudy (solomonzelman):

tell me what was your common ratio (r)? what was your number of terms? what is the first term?

OpenStudy (anonymous):

My common ratio was -5 The number of terms was 8 The first term was -3

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

\((1-(r^n))\) is what I meant

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle {\rm S}_8=\frac{ (-3)\times \left(1-(-5)^8\right)}{1-(-5)} }\) Click \(\bf\color{blue }{\href{http:///www.wolframalpha.com/input/?i=%28-3+%C3%97+%281+-+%28-5%29%5E8%29%29%2F%281-%28-5%29%29}{here}}\) to check the result.

OpenStudy (anonymous):

My common ratio was -5 The number of terms was 8 The first term was -3

OpenStudy (anonymous):

Is the answer 195,312??

OpenStudy (solomonzelman):

yes, that is the answer.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (solomonzelman):

anytime !